OpenStudy (anonymous):

completing the square x^2-8x+5

OpenStudy (shamil98):

x^2 - 8x + 5 = 0 add 11 to both sides x^2 - 8x + 16 = 11 (x-4)^2 = 11

OpenStudy (shamil98):

which becomes (x-4)^2 - 11